OpenStudy (poofypenguin):

If x is one of the sides of the rectangle with an area of 64 square meters, what is the domain of x?

OpenStudy (anonymous):

\[x \times x = 64^2\] \[x^2 = 64^2\] \[x=\sqrt{64}\] \[x=8\] \[8*8 = 8^2\]

OpenStudy (poofypenguin):

Umm... then what is the domain for x?

OpenStudy (anonymous):

@agentzero its not a square

OpenStudy (anonymous):

oops you are right my bad

OpenStudy (poofypenguin):

that's what i thought at first, but if the maximum is 64 doesn't that imply that the other side would be 1, but the other side could be a half, which would mean that x could be 128.

OpenStudy (anonymous):

hmm yes i have made an error, x is not necessarily an integer is it..

OpenStudy (anonymous):

 we just have 0<x